OpenStudy (anonymous):

how would you Graph f(x)= 2e^x-1?

OpenStudy (dumbcow):

do you know what the graph of f(x) = e^x looks like, its best to start there

OpenStudy (anonymous):

no i dont get wat i plug in for e

OpenStudy (dumbcow):

dont worry about the "e" its just a number. e = 2.72... you should have a button for it on your calculator more generally the graph of f(x) = b^x is a exponential function with a horizontal asymptote at y=0 and y-intercept at 1

OpenStudy (dumbcow):

btw is it e^(x-1) or (e^x) -1

OpenStudy (anonymous):

its (2e^x)-1

OpenStudy (dumbcow):

 Now shift the graph down 1 so new asymptote is y=-1 find new y-intercept y = 2e^0 -1 y = 2-1=1